I have tried solving for s in the equation but cannot seem to get it right--can you help?
5.00s +6.50s + 7.25s + 20 = 5.50s + 4.75s +5.75s +8.00s+ 18.0

If you could just get me pointed in the right direction I would appreciate this!

Answer by jim_thompson5910(35256)

Start with
5.00s +6.50s + 7.25s + 20 = 5.50s + 4.75s +5.75s +8.00s+ 18.00

and multiply EVERY number by 100 to clear out the decimals

100*5.00s +100*6.50s + 100*7.25s + 100*20 = 100*5.50s + 100*4.75s +100*5.75s +100*8.00s+ 100*18.00

500s +650s + 725s + 2000 = 550s + 475s +575s +800s+ 1800


Note: this effectively moves the decimal 2 spots to the right for each number

-------------------------------------------------------
Now let's solve for 's'


500s +650s + 725s + 2000 = 550s + 475s +575s +800s+ 1800

1875*s+2000 = 550s + 475s +575s +800s+ 1800

1875s+2000 = 2400s+1800

1875s - 2400s = 1800 - 2000

-525s = -200

s = -200/(-525)

s = 8/21

s = 0.3809523

If you need to round to the nearest hundredth, then the final answer is s = 0.38